OpenStudy (anonymous):

Solve the problem by stating the equation you would use to solve and then solving it. Be sure to show all of your work. After spending $3850 for tables and $1750 for chairs, a convention center manager finds that 30% of his original budget remains. Find the amount that remains. Round to the nearest dollar.

OpenStudy (anonymous):

x-x/3 = 3850+1750

OpenStudy (anonymous):

2x/3 = 5600

OpenStudy (anonymous):

x = 8400

OpenStudy (anonymous):

let the total amount before spending anything be = $x since 30% remains, spent is 70% = 3850 + 1750 so 70/100 * x = 5600 x = 5600 x 100/70 = 8000 So total in the beginning was 8000 left = total - spent = 8000 - 5600 =2400

OpenStudy (anonymous):

harkirat is more accurate mine is wrong sorry

OpenStudy (anonymous):

Hark is thee best bnb but thanx for ur input i really do enjoy and appreciate it